OpenStudy (calculusxy):

One pipe fills a pool in 4 hours. A second pipe, used to drain the water from the pool, can empty the pool in 8 hours. The owner of the pool mistakenly opened both pipes. How long will it take to fill the pool if one pipe is filling and the other is emptying?

OpenStudy (misty1212):

Hi!!

OpenStudy (calculusxy):

Hi!

OpenStudy (misty1212):

one has rate \(\frac{1}{4}\) the other has rate \(-\frac{1}{8}\) the combined rate is what you get when you ... combine them what is \[\frac{1}{4}-\frac{1}{8}\]?

OpenStudy (calculusxy):

\(\large \frac{2}{8} - \frac{1}{8} = \frac{1}{8}\)

OpenStudy (misty1212):

ok so the this system of silliness fills one eighths of the pool per hour

OpenStudy (misty1212):

how many hours to fill the whole pool?

OpenStudy (calculusxy):

i don't understand but i guess \(\large \frac{1}{8}\)?

OpenStudy (misty1212):

oh no

OpenStudy (misty1212):

it fills only one eighths of the pool per hour

OpenStudy (misty1212):

so how many hours for the entire pool to be filled?

OpenStudy (calculusxy):

okay so do we divide \(\large 1 \div \frac{1}{8}\) ?

OpenStudy (misty1212):

you can, or i would just think one eighths per hour, 8 hours for the whole pool

OpenStudy (calculusxy):

your thinking seems more logical. thank you!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]